Consider an economy that produces only three types of fruit: apples, oranges & bananas. In
the base year the production & price data are as follows:
FRUIT QUANTITY PRICE
APPLES 3000 Unit Rs. 2 per unit
BANANAS 6000 Rs. 3
ORANGES 8000 Rs. 4
In the current year the production & price data are as follows:
FRUIT QUANTITY PRICE
APPLES 4000 Unit Rs. 3 per unit
BANANAS 14,000 Rs. 2
ORANGES 32,000 Rs. 5
a. Find nominal GDP in the current year & in the base year. What is the percentage
increase since base year?
b. Find real GDP in the current year & in the base year. By what percentage does a real
GDP increase from the base year to current year?
c. Find the GDP deflator for the current year & the base year. By what percentage does
the price level change from the base year to current year?
1
Expert's answer
2016-11-04T12:11:10-0400
a. Nominal GDP in the base year is: Ynb = 3000*2 + 6000*3 + 8000*4 = Rs. 56,000.
Nominal GDP in the current year is: Ync = 4000*3 + 14,000*2 + 32,000*5 = Rs. 200,000.
The percentage increase since base year is (Ync/Ynb - 1)*100% = (200,000/56,000 - 1)*100% = 257.14%.
b. Real GDP in the base year is: Yrb = 3000*2 + 6000*3 + 8000*4 = Rs. 56,000.
Nominal GDP in the current year is: Yrc = 4000*2 + 14,000*3 + 32,000*4 = Rs. 178,000.
The percentage increase since base year is (Ync/Ynb - 1)*100% = (178,000/56,000 - 1)*100% = 217.86%.
c. The GDP deflator for the base year is GDP deflator = GDPnb/GDPrb*100 = 56,000/56,000*100 = 100.
The GDP deflator for the current year is GDP deflator = GDPnc/GDPrc*100 = 200,000/178,000*100 = 112.36.
The percentage change in the price level from the base year to current year is 112.36/100*100% = 12.36%.